Determine whether the graphs of each pair of lines are parallel.
3x+8=y
2y=6x-3
Are the graphs of the given equation parallel?
Thank you!

Answer by ewatrrr(24785)


Hi
Note:the standard slope-intercept form for an equation of a line is y = mx + b
where m is the slope and b the y-intercept.
3x+8=y OR y = 3x + 8 slope m = 3 PTs (0,8), (-8/3,0) ,(-2,2)
2y=6x-3 OR y = 3x -3/2 slope m = 3 PTs (0,-3/2), (1/2,0), (1,3/2))
Lines have identical slopes, therefore are parallel
Graphing